OpenStudy (anonymous):

Solve each system by elimination 6x+6y=-6 -x-6y=6

OpenStudy (anonymous):

Equation 1 : 6x+6y=-6 Equation 2 : -x-6y=6 multiply equation 2 by -1 Equation 2 now becomes x+6y=-6 Subtract the 2 equations 6x+6y=-6 -)x+6y=-6 ______________ 5x = 0 x=0 sub back into equation 1 6(0) +6y = -6 6y = -6 y = -1